Circular No. 7940

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002ed IN NGC 5468 B. Monard, Pretoria, South Africa, reports the discovery of an ap-

parent supernova (R mag 16.5) on unfiltered CCD frames taken with the Bronberg Observatory 0.31-m telescope on July 27.77 UT. The new object was confirmed on CCD frames taken on July 28.70 (R mag 16.4) by Monard. SN 2002ed is located at α = 14^{h}06^{m}38^{s}.2, δ = −5^{o}27′29′′ (equinox 2000.0), which is 55′′ east and 15′′ south of the center of NGC 5468. The object does not appear on CCD frames taken by Monard on July 2 and 23 (limiting R mag 17.5). W. Li, University of California at Berkeley, confirms the existence of the object on LOTOSS (cf. IAUC 7906) KAIT CCD images on July 24 to 29. NGC 5468 is the host galaxy for SN 1999cp (cf. IAUC 7205) and SN 2002cr (cf. IAUC 7890).

V2275 CYGNI C. C. Venturini, D. K. Lynch, R. J. Rudy, S. Mazuk, The Aerospace

Corporation; R. C. Puetter, Center for Astrophysics and Space Science, Univsersity of California at San Diego; and R. B. Perry, NASA Langley Research Center, report 0.8–2.5 μm spectroscopy using the Lick Observatory 3-m telescope and the Aerospace Corporation Near-Infrared Imaging Spectrograph on July 19.42 UT (∼ 335 days after discovery): “High-energy coronal lines dominate the spectrum with [Si X] at 1.4301 μm being the strongest. Other prominent coronal lines include [Si IX] at 1.2520 μm, [Al IX] at 2.0394 μm, and the rarely seen lines of [Fe XIII] at 1.0747 and 1.0798 μm. Many of the lower excitation coronal lines like [Ca VIII] and [Si VI] are weak or absent from the spectrum. He II lines and He I at 1.0830 μm are also present. There are large ranges in the emission line widths: the H I and He II lines have FWHM of approximately 900 km/s while the value for the [Si X] feature is 3400 km/s.”

SUPERNOVA 2002ap IN M74 K. Sárneczky, B. Sipőcz, Budapest, Hungary; and J. Vinkó, Szeged,

Hungary, report: “We have measured Johnson-Cousins CCD magnitudes with the 0.60-m Schmidt at Konkoly Observatory: July 27.05 UT, V = 17.12 ± 0.05; R = 16.24 ± 0.05; I = 16.05 ± 0.05. Compared with the data by Y. Yoshii et al. (cf. IAUC 7930), the light curve decline rates are 0.022, 0.019 and 0.024 mag/day in V, R and I bands, respectively.”

2002 July 29 \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT Carl W. Hergenrother